Harding, Jones, and Sandy are in the process of liquidating and the partners have the following capital balances; $24,000, $24,000, and ($9,000) respectively. The partners share all profits and losses 16%, 48%, and 36%, respectively. Sandy has indicated that the ($9,000) deficit will be covered with a forthcoming contribution. The remaining partners have requested to immediately receive $20,000 in cash that is available. How should this cash be distributed?
Answer #1

Hanson

James

Smith

24,000

24,000

-9000

16%

48%

36%



-9000

20,000






16/(16+48%)=25.00%

48/(16+48%)=75.00%


24,000 + 24,000

+ 20,000

=68,000



68,000*25%=17,000

68,000*75%=51,000



24,000+24,000=

48,000



51,000-48,000=3,000



answered by: Jack Who
Answer #2

Sandy has Indicated that the ($9,000) deficit will be covered with a forthcoming contribution

The Remaining Capital Balance is = 24000 + 24000 = 48000

Total cash Available = 20000

Loss = 48000 - 20000

Loss = $ 28000

Loss is to be shared between Harding& Jones in ratio = 16%:48%

Harding Capital balance thereafter = 24000 - 28000*16/(16+48) = $ 17000

Jones Capital balance thereafter = 24000 - 28000*48/(16+48) = $ 3000

Cash to be Distributed in their capital balance ratio

Therefore Cash to be distributed to

Harding = $ 17000

Jones = $ 3000

Answer

Harding $ 17000 ; Jones $ 3000